lolokrat:

Can someone check my geometry homework? I’ve been sick so I missed a few lectures on this so was unsure how to do this. Some of them I probably bombed but if anyone could help it’d be much appreciated!

silvernight269:

i would say side 1 #4 would be supplementary.

silvernight269:

oh wait nvm ur right

silvernight269:

13 said to name a linear PAIR and 14 would be angle CGA

lolokrat:

how do i write a linear pair? sorry never learned this haha

silvernight269:

they meant a pair of angles

lolokrat:

<EGD , <CGD would this be a linear pair?

silvernight269:

no

silvernight269:

linear means the same <EGC is coLINEAR meaning its linear but on the same line <FGC and < CGD there are others but if you want to use that then thats fine too

lolokrat:

so fgc and cgd would be a linear pair?

silvernight269:

yes

lolokrat:

ok. is there anything else i missed on that page?

silvernight269:

nope now the second pg #2 second angle. its <HKG not <HKJ

lolokrat:

ohh yeah

silvernight269:

5 and 6 are also wrong too.

lolokrat:

5 would it be EKJ ?

lolokrat:

#6* sorry

silvernight269:

For #5, two angles have to equal 90 degree. the ones u put down are supplementary.

lolokrat:

#5 FKG and GKJ would this be it?

silvernight269:

and you guessed wrong. its <GKJ for #5

silvernight269:

there is no second angle for #5

silvernight269:

the answer would be <FKE for #6

lolokrat:

isnt <EKF and <FKE this same thing?

silvernight269:

yea they both are the same so it doesnt matter which one to choose from, just one of those 2 angles.

lolokrat:

oh?? I had put that down the first time and you said it was wrong??

silvernight269:

oh yea oops sorry. didnt see it backwards

lolokrat:

oh okay good i was confused lol

silvernight269:

anyways, for #13, i dont see the m<

lolokrat:

oh yeah i got confused and moved onto the next

silvernight269:

calculator? and for number 15

lolokrat:

when adding them up should they be the same for 13 i got 4(24)+4=100 3(24)+1=73 they arent the same??

silvernight269:

thats strange

silvernight269:

u calculated the first equation wrong its 4(24)+11=107

lolokrat:

ohh i did +4 for some reason

silvernight269:

yea at first i thought so too but looking back at the TYPED equations i figured it was 11. for #17, if the answer 90 degree for all angles? including angle 3?

silvernight269:

and the rest is pretty much correct! Great Job!

lolokrat:

wait what about #17??? how would i do it?

lolokrat:

and for #15 i put the measures 38+10=48 3(38)+18=132

silvernight269:

ok good

lolokrat:

so #17 how would it work??

silvernight269:

<3 is a 90 degree but you need to write <4 and <5 down

silvernight269:

<3 is 90 degrees because <4 and <5 added up equal to 90 degree meaning that the angles around it all have to add up to 360 degree.

silvernight269:

Get it?

lolokrat:

er still not understanding

lolokrat:

what did i write down?? are those not angles 4 and 5??

silvernight269:

90 degree is not angles 4 and 5. it is the sum of both of those angles

lolokrat:

can you work it out so i can see what u mean?? im totally lost right now

silvernight269:

circle m<4= 31 m<5=51

silvernight269:

59* my bad

lolokrat:

m<4=2(18)-5=31 m<5=4(18)-13=59

lolokrat:

this?

silvernight269:

yea

lolokrat:

oh so did i get it right and just not label or??

silvernight269:

you didnt circle it meaning u only circled 90

lolokrat:

circled?

silvernight269:

nvm. other term is you didnt label it.

lolokrat:

ohh ok lol

silvernight269:

lmao